OpenStudy (anonymous):

2.5kg of dressed chicken cost 105. what will be the cost of 5.75?

OpenStudy (anonymous):

i mean 5.75 kg

jhonyy9 (jhonyy9):

so 2,5 --- 105 5,75 --- x -------------- x= you need calcule the x from this system ok ?

jhonyy9 (jhonyy9):

so this is the rule of 3 simples

jhonyy9 (jhonyy9):

can you doit ?

OpenStudy (anonymous):

no .. pls help me...

jhonyy9 (jhonyy9):

just cross multiplie and divide both sides by 2,5 and will get the value of x

jhonyy9 (jhonyy9):

2,5x=105*5,75 divide both sides by 2,5 x= (105*5,75)/2,5 = hope you can calcule it sure easy

OpenStudy (anonymous):

thanks .. :))

jhonyy9 (jhonyy9):

was my pleasure good luck bye